# A case for specialisation in non-human entities

### v1

@@ Abstract
Our contribution is threefold. First, we review the most widely accepted arguments against specialisation, and discuss how their relevance in the context of human labour is actually an argument for specialisation in the case of non human agents, be they algorithms or human organisations. Second, we propose four arguments in favor of specialisation, ranging from machine learning robustness, to computer security, social sciences and cultural evolution. Third, we finally make a case for specification, discuss how the machine learning approach to AI has so far failed to catch up with good practices from safety-engineering and formal verification of software, and discuss how some emerging good practices in machine learning help reduce this gap. In particular, we justify the need for specified governance for hard-to-specify systems.
@@ 1 Introduction
In 2023, the European Union finally announced its long-awaited Artificial Intelligence Act (AI Act). The AI act consists in a series of regulations on AI that not only impacts European AI, but also ultimately influences other countries and regions in the world in what is known as the Brussels’ effect[Bra20, Bra24, SA22, GDV21]. One of the most puzzling features of the AI Act is that, due to intensive lobbying from large AI models companies[Per23], the Act ended up putting more restrictions on some specialised AI models than it does on those claiming general capabilities. To capture the absurdity of the situation, we propose the following analogy: can we imagine if in medical practice, a general practitioner had more freedom to operate on a patient’s eye than an ophthalmologist? The answer is obviously negative. The less specialised a medical doctor, the more restrictions different medical legislation or codes of ethics[Ass71, dls53] would impose on what they can do.
@@ 2 Context
Historically, the word “algorithm” comes from the name of Muhammad ibn Musa al-Khwarizmi, and really consisted of decomposing complex tasks into elementary trivial subtasks. This intuitive approach was further formalized by the pioneers of computer science, in particular Alonzo Church [Chu36] and Alan Turing [Tur36]. Both argued that any computable function can be written as a composition of elementary ultra-specialised hard-wired mechanical operations, typically on a Turing machine (or in lambda calculus). In today’s computing machines, such operations are the (now numerous) logical circuits engraved in processing units like CPUs. The assertion that these fixed circuits suffice to perform any computable information processing function is now known as the Church-Turing thesis. This thesis arguably thus asserts that, in principle, any general capability is a composition of (a very large number of) specialised operations.
@@
In fact, many software developers might not view their software solutions as a composition of specialised operations, especially if their solutions leverage external libraries. While some of these codes simply translate the developers’ programming language codes into elementary binary codes (these are compilers), it has become extremely common for developers to use complex “general-purpose” libraries, especially when importing, e.g., large language models. Such systems could be argued to “empower” developers, as they can now effortlessly program solutions to many more problems. However, they can also be argued to reduce their capability to understand their own systems, especially in terms of capability, safe usage and emergent risks.
What we might call “specialisation blindness” can be found in other activities or professions. A fast-food franchisee, for example, considers that their business is selling hamburgers or other products sold in this type of restaurant. However, upstream, a franchisor has to call on several specialist trades to develop all the building blocks for a successful franchise: the right furniture (interior designers, furniture salesmen…), the right recipes (chefs, nutritionists, etc.), the right ingredients (farms, bakers, drinks wholesalers…), the right business model (bankers, accountants…), the right location (estate agents, notaries…), etc.
Specialisation blindness is arguably an unavoidable consequence of collaborating to perform complex tasks. However, the question we raise in this paper, is how to organise collaboration to make it an asset rather than a risk. Essentially, we will argue that the key lies in well-specified specialisation.
2.1 Defining generality
In the context of the AI Act, generality is often associated with the purpose of a system. The greater the diversity of use cases, the more “general-purpose” the system. This definition of generality-by-purpose can be also found in Model Cards [MWZ+19], which asks AI system developers to specify the “intended use cases” of the developed systems.
@@ 2.1 Defining generality
A leaked document [Max24] from OpenAI and Microsoft revealed their use of a more financial approach to defining AGI. Namely, they (privately) regarded as “general” a system that generates hundreds of billions of dollars in revenues. In a sense, this approach is similar to counting use cases of a system, but it proposes to weigh the use cases by their financial added values. It is however noteworthy that, given this definition, Blackrock’s trading algorithm Aladdin  [^2] In 2020, Aladdin managed 21.6 trillion U.S. dollars in assets[Ung20]., and probably also the algorithms of Google’s AdSense  [^3] Google’s ad revenue amounted to 237.86 billion U.S. dollars in 2023[Bia24]. should be regarded as much more “general” than language models.
@@
Note that a very specialised system in this sense may nevertheless be extremely complex. For instance, even though a language model is only accessible through a “next-token prediction” API, or if its only API yields content recommendations, it may be itself composed of trillions of parameters [LYZ+22]. In fact, any large human organization that is specialised to deliver a very specific task must perform a large number of internal tasks, such as accounting, sales or human resource management. Nevertheless, we will regard it as specialised if the number of services it offers to external users is small (and well-specified).
2.2 Defining the notion of task and its granularity
@@
When we look at the history of the organisation and rationalisation of work, particularly through the emblematic examples of Taylorism [Lit78], Fordism [Wat19] and Toyotism [DJN85], all of these paradigms have involved defining and sorting out specific tasks, selecting workers to carry them out, and more generally (over)specialising work and then automating it. For the past two centuries, automation has been the hallmark of work transformation. Broadly speaking, it starts with an agent’s craft know-how. Automation then consists of extracting the skills from the agent, and of encoding them into a procedure, i.e. a series of simple, structured and repetitive tasks be they physical or cognitive. In debates about the “future of work” [BM17], jobs that are identified as least likely to be automated are those whose tasks are unstructured, non-routine, and whose performance is associated with critical thinking, long chains of reasoning or complex planning, a certain level of creativity, etc. Arguably, this is because the resolutions of such tasks are hard to encode into procedures, though the rise of machine learning opens the door to the encoding of procedures that are humanly hard to describe, e.g. in the Kolomogorov-Solomonoff sense (see Section 5.4).
@@
Automation - the creation of machines and algorithms to carry out tasks previously performed by humans - is a constant reconsideration of division of labour. To quote Karl Marx (himself echoing Charles Babbage), machines emerge as a synthesis of the division of labour [Mar47]. So the notion of task is at the heart of automation. An important question is to know what level of definition and operations to adopt in order to automate tasks and, ultimately, specialisation. There are limitations that make this exercise complicated. A prominent challenge is to select the granularity of a task (and thus its level of specialisation). There are informational and computational hurdles that make this exercise complicated. For example, there are many tasks that we understand tacitly and carry out without being able to state their explicit rules and procedures. In other words, formulated by Polanyi’s paradox [Aut14], “we can know more than we can tell” [Pol09]. Moreover, automating certain tasks and not others also means renouncing possible ways of working and other possible uses [Sim58].
@@
Defining a task can also refer to expressions such as granularity or modularity. We can define granularity and modularity using the definitions proposed by Benkler [Ben06] which are the basis of his socio-economic production model “Commons-based peer production”. Modularity is a way of dividing a project into smaller components (modules) that can be developed independently before being assembled to form a whole, in order to maximise the flexibility and autonomy of the contributors. Granularity refers to the size of the modules, defined in terms of the time and energy required by the participants to produce them. It is important to specify the modules and their quantity in a project, because the interest and investment of the participants depend on it. As we will see, the two concepts are also central in computer science and system engineering.
@@ 3.1 The case of humans
Although some arguments may have been anterior, the most severe criticisms of the division of labor opposed Taylorism, assembly line work, and what came to be known as Scientific Management[BD18]. This work organisation generated an extreme division of labour with a fragmentation of tasks, a sustained pace of execution, a high degree of dependence between workers, and the impossibility of envisaging the unexpected events. Criticisms from the social sciences of work were to emerge in the middle of the twentieth century and the rise of increasing industrialisation (particularly in the automobile sector) and mechanisation. Friedmann [Fri55, Fri61] talks of “work in crumbs” [^4] Better captured by the original French expression “Travail en miettes”.” to describe the evolution of forms of work with a high level of automation, and in particular the work in the factories. There have also been concerned with the psychological damage caused to workers by alienating, unskilled, and depersonalized work.
Today, criticisms of the excessive division of labor and hyper-specialisation concern many economic sectors (food industry, logistics, textile industry, etc.). They particularly focus on the aspects that are detrimental to the body and mind of individuals. One of these industries is precisely the information and communications technology industry, and particularly digital platforms (on-demand platforms, social media), and more generally artificial intelligence using digital labor, micro-work and data workers, especially in the global south countries [Cas20, WMG22].
@@
High specialisation can cause workers to lose interest and motivation in their work, and can impact performance [LLCD09]. A high level of specialisation tends to lower workers’ stimulation and motivation levels, while increasing their boredom and disengagement [Hac69, MROM94].
@@
However, while the psychological and physical damages on human workers are important to consider, algorithms are arguably not subject to such concerns. At least following [GM22], while they could be regarded as moral agents (i.e. they have moral duties), they should not be regarded as moral patients (i.e. there is no moral duty to protect them). Therefore, the above arguments do not seem to apply to non-human agents.
3.2 Arguments from economics
@@
In economics, instead of specialising in a given step of the production chain, companies may be tempted to invest in vertical integration [Coa37, BL12]. Namely, they could want to own their entire supply and delivery chain. The theoretical analysis of [Arr75] indeed found that imperfect information can incentivize such vertical integration, while that of [Car79] instead stresses the role of transaction costs. [GH86] further underlined contractualisation costs: if specifying the conditions of an agreement between a supplier and a producer is challenging, then there will be strong incentives for vertical integration. The theories of [Arr75] and [Car79] have both found support, e.g. by [Lie91] in the chemical industry. On top of this, [FF10] empirically observed better investment coordination and less financial risk following the vertical integration of the Swiss electricity sector. In practice, varying working conditions may instead incentivize outsourcing, and thus vertical disintegration [Ric17, KK05].
@@
In the software industry, yet another phenomenon may incentivize both vertical and lateral integration, namely the network effect [Sha99]. This describes situations where the value of a product is increased by the wide use of this product, or of similar related products. A classical instance of this is Facebook’s huge investments in social media. For instance, by acquiring Instagram and by connecting it with their other platforms, Facebook has exploited the network effect to increase the value of their anterior assets [LA17]. Similarly, Google and Apple now produce both hardware, operating systems and applications, thereby offering a unifying solution to their customers. The value of the hardware is then augmented by a corresponding optimized optimized system, which is itself augmented by its connection to a dedicated cloud system [VDR12]. Further, software companies have incentives to invest in lateral integration, to propose all-in-one highly connected digital workspaces, as is the case for instance of the Microsoft 365 product. This further helped them push AI-based solutions, such as Microsoft co-pilot [SK12].
@@
However, while there may be some positive consequences for some users, critics point out that such integrations incur societal risks, especially in terms of market power [LP97]. In particular, the technology sector has been repeatedly criticized for its dependence on a small number of actors, as exemplified by the landmark antitrust case against Microsoft [Eco01], and by the numerous recent calls to enforce antitrust laws to other actors [MKSS24, DG24, Wör24].
@@
Additionally, integration can increase systemic risks. This is well illustrated by the report of the Cybersecurity and Infrastructure Security Agency on the 2023 Microsoft Online Exchange Incident [SA24]. The penetration of Microsoft’s cloud infrastructure by foreign actors has not only compromised the targeted US institutions; it may also be endangering all of Microsoft’s customers. Moreover, the high connectivity of all of Microsoft’s cloud services means that a given customer may be compromised, even if they only use some of these services.
@@
Having said this, at the scale of a country, especially in a context of tensed geopolitical tension, vertical integration could in fact be a consideration of national security. More generally, it could be in a given system’s best interest to seek generality, if they want autonomy and resilience. However, it may not be other others’ best interest to depend on a general system. In any case, we recall that our case for specialisation is focused on external uses of a system.
3.3 Arguments from statistics
@@
In artificial intelligence, much of the recent interest for general AIs derives from the observed “scaling laws” [KMH+20a], which argue that learning from all sorts of (non-specialised) data is the driving force of dramatic performances. While these observations have been criticized [DM24], they do have some theoretical backgrounds.
@@
In particular, they may be argued to be an instance of Stein’s paradox [EM77]. Strikingly, this paradox shows that, when performing statistical learning on three or more disjoint subsets of data, it is statistically inadmissible to learn separately on each subset. More precisely, for each subset, consider any (specialised) estimator of a ground truth statistics of the subset. Then there is a (general) estimator that learns from the union of all subsets such that, no matter what the ground truth is, the general estimator’s expected mean square error will be lower than the specialised estimators’. Moreover, for at least one value of the ground truth, it is strictly lower. The general estimator is said to strictly dominate the specialised estimators.
@@
This is especially evident in the case of collaborative filtering [ERK+11]. To better optimize what content should be recommended to a given user, it is extremely useful to leverage the preferences of similar users, rather than to learn exclusively from the given user’s data.
@@
But more importantly, the subsets of data have thus far been assumed to safe to learn from. However, in practice, most datasets raise privacy and poisoning issues. On one hand, general estimator can more easily cross information that would allow for user de-anonymisation, even if “privacy-preserving learning” solutions like differential privacy are used [KM11, ZXLZ15]. On the other hand, malicious users can more easily poison the union of all subsets, thereby biasing or harming the general estimator [BNL12, SMS+21, FGHV22]. In particular, the introduction of collaborative filtering has given fake accounts an enormous influence on the daily information exposure of billions of humans.

### v2

@@ Abstract
Our contribution is threefold. First, we review the most widely accepted arguments against specialisation and discuss how their relevance in the context of human labour is actually an argument for specialisation in the case of non human agents, be they algorithms or human organisations. Second, we propose four arguments in favor of specialisation, ranging from machine learning robustness, to computer security, social sciences and cultural evolution. Third, we finally make a case for specification, discuss how the machine learning approach to AI has so far failed to catch up with good practices from safety-engineering and formal verification of software, and discuss how some emerging good practices in machine learning help reduce this gap. In particular, we justify the need for specified governance for hard-to-specify systems.
@@ 1 Introduction
In 2023, the European Union announced its Artificial Intelligence Act. The AI Act consists in a series of regulations on AI that not only impacts European AI, but ultimately also other countries and regions in the world in what is known as the Brussels’ effect (Bradford 2020, 2024; Siegmann and Anderljung 2022; Gunst and De Ville 2021). One of the most puzzling features of the AI Act is that, due to intensive lobbying from large AI companies (Perrigo 2023), the Act ended up putting more restrictions on some specialised AI models than it does on AI models claiming general capabilities. To capture the absurdity, consider an analogy: in medical practice, would a general practitioner have more freedom to operate on a patient’s eye than an ophthalmologist? The answer is obviously negative. The less specialised a medical doctor, the more restrictions different medical legislation or codes of ethics [^1] See e.g.,  (American Medical Association 1871; codesante 2004 (initial version 1953Ministère de la santé)). would impose on what they can do.
@@ 2 Context
Historically, the word “algorithm” comes from the name of Muhammad ibn Musa al-Khwarizmi, and really consisted of decomposing complex tasks into elementary trivial subtasks, just like in al-Khawarizmi’s pioneering book, providing simple step-by-step recipes to simplify problem resolution for a wider audience, and in particular, inheritance case resolution for lawyers.
This intuitive approach, insisting on breaking problem resolution to simple tasks, was further formalized by the pioneers of computer science, in particular Alonzo Church (Church 1936) and Alan Turing (Turing 1936). Both argued that any computable function can be written as a composition of elementary ultra-specialised hard-wired mechanical operations, typically on a Turing machine (or in lambda calculus). In today’s computing machines, such operations are the (now numerous) logical circuits engraved in processing units like CPUs. The assertion that these fixed circuits suffice to perform any computable information processing function is now known as the Church-Turing thesis. This thesis arguably thus asserts that, in principle, any general capability is a composition of (a very large number of) specialised operations.
@@
In fact, many software developers might not view their software solutions as a composition of specialised operations, especially if their solutions leverage external libraries. While some of these codes simply translate the developers’ programming language codes into elementary binary codes (these are compilers), it has become extremely common for developers to use complex “general-purpose” libraries, especially when importing, e.g., large language models. Such systems could be argued to “empower” developers, as they can now effortlessly program solutions to many more problems. However, they can also be argued to reduce their capability to understand their own systems, especially in terms of capability, safe usage and emergent risks. In contrast, developers could favor building upon specialised libraries, which provide them with precisely the tools they needed for some well-defined tasks, such as optimization solvers. The developer would then only be blind to such specialised operations, which potentially having the capability to clean and sanitize their outputs. It is noteworthy that such specialised libraries could be made to be equally blind to how they are used.
What we might call “specialisation blindness” can also be found in other activities or professions. A fast-food franchisee, for example, considers that their business is selling hamburgers or other products sold in this type of restaurant. However, upstream, a franchisor has to call on several specialist trades to develop all the building blocks for a successful franchise: the right furniture (interior designers, furniture salesmen…), the right recipes (chefs, nutritionists, etc.), the right ingredients (farms, bakers, drinks wholesalers…), the right business model (bankers, accountants…), the right location (estate agents, notaries…), etc.
Specialisation blindness is arguably an unavoidable consequence of collaborating to perform complex tasks (Boullier and El Mhamdi 2020; Tighanimine 2025). However, the question we raise in this paper, is how to organise collaboration to make it an asset rather than a risk. Essentially, we will argue that the key lies in well-specified specialisation.
Defining Generality
In the context of the AI Act, generality is often associated with the purpose of a system. The greater the diversity of use cases, the more “general-purpose” the system. This definition of generality-by-purpose can be also found in Model Cards (Mitchell et al. 2019), which asks AI system developers to specify the “intended use cases” of the developed systems.
@@
A leaked document (Maxwell 2024) from OpenAI and Microsoft revealed their use of a more financial approach to defining AGI. Namely, they (privately) regarded as “general” a system that generates hundreds of billions of dollars in revenues. In a sense, this approach is similar to counting use cases of a system, but it proposes to weigh the use cases by their financial added values. It is however noteworthy that, given this definition, Blackrock’s trading algorithm Aladdin  [^2] In 2020, Aladdin managed 21.6 trillion U.S. dollars in assets (Ungarino 2020). and Google’s ad targeting algorithm AdSense  [^3] Google’s ad revenue amounted to 237.86 billion U.S. dollars in 2023 (Bianchi 2024). should be regarded as much more “general” than language models.
@@
Note that a very specialised system in this sense may nevertheless be extremely complex. For instance, even though a language model is only accessible through a “next-token prediction” API, or if its only API yields content recommendations, it may be itself composed of trillions of parameters (Lian et al. 2022). In fact, any large human organization that is specialised to deliver a very specific task must perform a large number of internal tasks, such as accounting, sales or human resource management. Nevertheless, we will regard it as specialised if the number of services it offers to external users is small (and well-specified).
AI Agents
Since 2024, most large AI companies have been promoting AIs with acting capabilities (Yao et al. 2023; Acharya, Kuppan, and Divya 2025; Murugesan 2025; Borghoff, Bottoni, and Pareschi 2025). As opposed to conversational AIs that merely send messages to end users, such so-called AI agents (or agentic AI) can execute commands on information systems. More precisely, the Model Context Protocol (MCP) (Hou et al. 2025; Ray 2025) was recently proposed to standardise the way AI agents do so. Namely, each AI agent is given a list of API calls that they can call, along with the documentation of these API end points.
It is noteworthy that such agents differ from other concepts of autonomous agents that rather refer to the capability of such agents to behave well in some general frameworks, like reinforcement learning (Hutter 2003; Veness et al. 2010), often with respect to a given objective function that sends rewards (Ben-Porat et al. 2024; Kierans et al. 2025). Such frameworks typically do not invoke the access to API calls.
In some regards, MCP thereby somewhat formalizes the concept of generality we discussed earlier. In particular, one could be tempted to give to an AI agent a large number of API call accesses, to enable them to solve complex tasks. Our case is especially against such developments. In particular, like others (Blili-Hamelin et al. 2025; Radosevich and Halloran 2025), we argue that it is not in the interest of companies and societies to integrate such agents in their organizations.
Defining the Notion of Task and its Granularity
@@
When we look at the history of the organisation and rationalisation of work, particularly through the emblematic examples of Taylorism (Littler 1978), Fordism (Watson 2019) and Toyotism (Dohse, Jürgens, and Nialsch 1985), all of these paradigms have involved defining and sorting out specific tasks, selecting workers to carry them out, and more generally (over)specialising work and then automating it. For the past two centuries, automation has been the hallmark of work transformation. Broadly speaking, it starts with an agent’s craft know-how. Automation then consists of extracting the skills from the agent, and of encoding them into a procedure, i.e. a series of simple, structured and repetitive tasks be they physical or cognitive. In debates about the “future of work” (Brynjolfsson and Mitchell 2017), jobs that are identified as least likely to be automated are those whose tasks are unstructured, non-routine, and whose performance is associated with critical thinking, long chains of reasoning or complex planning, a certain level of creativity, etc. Arguably, this is because the resolutions of such tasks are hard to encode into procedures, though the rise of machine learning opens the door to the encoding of procedures that are humanly hard to describe, e.g. in the Kolomogorov-Solomonoff sense (see Section 5).
@@
Automation - the creation of machines and algorithms to carry out tasks previously performed by humans - is a constant reconsideration of division of labour. To quote Karl Marx (himself echoing Charles Babbage), machines emerge as a synthesis of the division of labour (Marx 1847). So the notion of task is at the heart of automation. An important question is to know what level of definition and operations to adopt in order to automate tasks and, ultimately, specialisation. There are limitations that make this exercise complicated. A prominent challenge is to select the granularity of a task (and thus its level of specialisation). There are informational and computational hurdles that make this exercise complicated. For example, there are many tasks that we understand tacitly and carry out without being able to state their explicit rules and procedures. In other words, formulated by Polanyi’s paradox (Autor 2014), “we can know more than we can tell” (Polanyi 2009). Moreover, automating certain tasks and not others also means renouncing possible ways of working and other possible uses (Simondon 1958).
@@
Defining a task can also refer to expressions such as granularity or modularity. We can define granularity and modularity using the definitions proposed by Benkler (Benkler 2006) which are the basis of his socio-economic production model “Commons-based peer production”. Modularity is a way of dividing a project into smaller components (modules) that can be developed independently before being assembled to form a whole, in order to maximise the flexibility and autonomy of the contributors. Granularity refers to the size of the modules, defined in terms of the time and energy required by the participants to produce them. It is important to specify the modules and their quantity in a project, because the interest and investment of the participants depend on it. As we will see, the two concepts are also central in computer science and system engineering.
@@ 3 The Case Against Specialisation
Although some arguments may have been anterior, the most severe criticisms of the division of labor opposed Taylorism, assembly line work, and what came to be known as Scientific Management(Bookwalter Drury 1918). This work organisation generated an extreme division of labour with a fragmentation of tasks, a sustained pace of execution, a high degree of dependence between workers, and the impossibility of envisaging the unexpected events. Criticisms from the social sciences of work were to emerge in the middle of the twentieth century and the rise of increasing industrialisation (particularly in the automobile sector) and mechanisation. Friedmann (Friedmann 1955; Friedman 1961) talks of “work in crumbs” [^4] Better captured by the original French expression “Travail en miettes”.” to describe the evolution of forms of work with a high level of automation, and in particular the work in the factories. There have also been concerned with the psychological damage caused to workers by alienating, unskilled, and depersonalized work.
Today, criticisms of the excessive division of labor and hyper-specialisation concern many economic sectors (food industry, logistics, textile industry, etc.). They particularly focus on the aspects that are detrimental to the body and mind of individuals. One of these industries is precisely the information and communications technology industry, and particularly digital platforms (on-demand platforms, social media), and more generally artificial intelligence using digital labor, micro-work and data workers, especially in the global south countries (Casilli 2020; Williams, Miceli, and Gebru 2022).
@@
High specialisation can cause workers to lose interest and motivation in their work, and can impact performance (Loukidou, Loan-Clarke, and Daniels 2009). A high level of specialisation tends to lower workers’ stimulation and motivation levels, while increasing their boredom and disengagement (Hackman 1969; McCauley et al. 1994).
@@
However, while the psychological and physical damages on human workers are important to consider, algorithms are arguably not subject to such concerns. At least following (Gibert and Martin 2022), while they could be regarded as moral agents (i.e. they have moral duties), they should not be regarded as moral patients (i.e. there is no moral duty to protect them). Therefore, the above arguments do not seem to apply to non-human agents.
Arguments from Economics
@@
In economics, instead of specialising in a given step of the production chain, companies may be tempted to invest in vertical integration (Coase 1937; Bresnahan and Levin 2012). Namely, they could want to own their entire supply and delivery chain. The theoretical analysis of (Arrow 1975) indeed found that imperfect information can incentivize such vertical integration, while that of (Carlton 1979) instead stresses the role of transaction costs. (Grossman and Hart 1986) further underlined contractualisation costs: if specifying the conditions of an agreement between a supplier and a producer is challenging, then there will be strong incentives for vertical integration. The theories of (Arrow 1975) and (Carlton 1979) have both found support, e.g. by (Lieberman 1991) in the chemical industry. On top of this, (Fetz and Filippini 2010) empirically observed better investment coordination and less financial risk following the vertical integration of the Swiss electricity sector. In practice, varying working conditions may instead incentivize outsourcing, and thus vertical disintegration (Ricardo 1817; Kakabadse and Kakabadse 2005).
@@
In the software industry, yet another phenomenon may incentivize both vertical and lateral integration, namely the network effect (Shapiro 1999). This describes situations where the value of a product is increased by the wide use of this product, or of similar related products. A classical instance of this is Facebook’s huge investments in social media. For instance, by acquiring Instagram and by connecting it with their other platforms, Facebook has exploited the network effect to increase the value of their anterior assets (Li and Agarwal 2017). Similarly, Google and Apple now produce both hardware, operating systems and applications, thereby offering a unifying solution to their customers. The value of the hardware is then augmented by a corresponding optimized optimized system, which is itself augmented by its connection to a dedicated cloud system (Vergara and Del Rosario 2012). Further, software companies have incentives to invest in lateral integration, to propose all-in-one highly connected digital workspaces, as is the case for instance of the Microsoft 365 product. This further helped them push AI-based solutions, such as Microsoft co-pilot (Skendzic and Kovacic 2012).
@@
However, while there may be some positive consequences for some users, critics point out that such integration incur societal risks, especially in terms of market power (Landes and Posner 1997). In particular, the technology sector has been repeatedly criticized for its dependence on a small number of actors, as exemplified by the landmark antitrust case against Microsoft (Economides 2001) and the current ongoing cases against Google and Facebook (Brennan 2025), and by the numerous recent calls to enforce antitrust laws to other actors (Munir et al. 2024; Davies and Georgieva 2024; Wörsdörfer 2024).
@@
Additionally, integration can increase systemic risks. This is well illustrated by the report of the Cybersecurity and Infrastructure Security Agency on the 2023 Microsoft Online Exchange Incident (Silvers and Alperovitch 2024). The penetration of Microsoft’s cloud infrastructure by foreign actors has not only compromised the targeted US institutions; it may also be endangering all of Microsoft’s customers. Moreover, the high connectivity of all of Microsoft’s cloud services means that a given customer may be compromised, even if they only use some of these services.
@@
Having said this, at the scale of a country, especially in a context of tensed geopolitical tension, vertical integration could in fact be a consideration of national security. More generally, it could be in a given system’s best interest to seek generality, if they want autonomy and resilience. However, it may not be other others’ best interest to depend on a general system, which is the consideration we will focus on. Indeed, we recall that our case for specialisation is focused on external uses of a system.
Arguments from Statistics
@@
In artificial intelligence, much of the recent interest for general AIs derives from the observed “scaling laws” (Kaplan et al. 2020a), which argue that learning from all sorts of (non-specialised) data is the driving force of dramatic performances. While these observations have been criticized (Diaz and Madaio 2024), they do have some theoretical backgrounds.
@@
In particular, they may be argued to be an instance of Stein’s paradox (Efron and Morris 1977). Strikingly, this paradox shows that, when performing statistical learning on three or more disjoint subsets of data, it is statistically inadmissible to learn separately on each subset. More precisely, for each subset, consider any (specialised) estimator of a ground truth statistics of the subset. Then there is a (general) estimator that learns from the union of all subsets such that, no matter what the ground truth is, the general estimator’s expected mean square error will be lower than the specialised estimators’. Moreover, for at least one value of the ground truth, it is strictly lower. The general estimator is said to strictly dominate the specialised estimators.
@@
This is especially evident in the case of collaborative filtering (Ekstrand et al. 2011). To better optimize what content should be recommended to a given user, it is extremely useful to leverage the preferences of similar users, rather than to learn exclusively from the given user’s data.
@@
But more importantly, the subsets of data have thus far been assumed to safe to learn from. However, in practice, most datasets raise privacy and poisoning issues. On one hand, general estimator can more easily cross information that would allow for user de-anonymisation, even if “privacy-preserving learning” solutions like differential privacy are used (Kifer and Machanavajjhala 2011; Zhu et al. 2015). On the other hand, malicious users can more easily poison the union of all subsets, thereby biasing or harming the general estimator (Biggio, Nelson, and Laskov 2012; Suya et al. 2021; Farhadkhani et al. 2022). In particular, the introduction of collaborative filtering has given fake accounts an enormous influence on the daily information exposure of billions of humans.